OpenStudy (anonymous):

there are 49 students in ms. coffmen's 3'ed period class. of those students 3/7 of them are male. how many females are in the class

OpenStudy (unklerhaukus):

4/7 x 49 =28

OpenStudy (anonymous):

(3/7) x 49= 21 males 49- 21 males= 28 males

OpenStudy (anonymous):

Lol, i meant 28 hoes...

OpenStudy (anonymous):

but how did you get the multipiclation for 49 times 3/7

OpenStudy (unklerhaukus):

know your times tables 49 is 7x7

OpenStudy (anonymous):

because (3/7)= .428571 x 100= 42.8571% of male in the class. So if you take the total # of students times the that percentage, you'll get the number of male. 49 students x 42.8571%= 21 gay dudes If 21 of the 49 are gay, that makes the other 28 shemales...